# Foundation Models for Structural Health Monitoring

### v1

@@ 6.4 Impact of pre-training
The test MAEobtained by our foundation model in the three cases are reported in Figure 8. As shown, No Pretrain always performs worse than the other two setups, underlying the usefulness of self-supervised pre-training. Quantitatively, Pretrain All reduces the MAEby 1.94$\times$, 3$\times$ and 1.2$\times$ on UC2-Light, UC2-Heavy and UC3, respectively, compared to No Pretrain. As expected, the benefits of pre-training are more evident for UC2, which has the smallest dataset. Moreover, pre-training the foundation model on all three datasets shows the best results for UC3 (with Pretrain All achieving 1.2$\times$ lower MAEcompared to Pretrain UC), while no clear advantage is evident for UC2. This is in accordance with the observation of the previous section that UC3 is the most difficult downstream tasks among the three, thus benefitting more from high-quality and generalizable representations. These findings indicate that the effectiveness of this approach might be further amplified by leveraging substantially larger and more diverse datasets for pre-training.
@@ 6.5 Model size optimization and Knowledge Distillation
Figure 9: MAEcomparison between regular and KD-enhanced fine-tuning.
We analyze the trade-off between model size and performance. For all UCs, following the findings of the previous section, we use a Pretrain All scheme, and we compare how regular supervised fine-tuning stands against KD-enhanced fine-tuning. We focus on TLE tasks because, for UC1, we were able to obtain the same performance shown in Figure 5 for the largest transformer, with a model of just $0.2$MB, with $(e_{dim},d_{dim})=(24,16)$. Results for UC2 and UC3 are shown in Figure 9. For clarity’s sake, we present results only in terms of MAE%, as other metrics follow similar trends. As shown, KD helps improve the performance of the models in almost all configurations, with the exception of a few cases related to UC2, especially in the low-size regime, where it provides marginal advantages. This is motivated by the fact that student networks that are too small cannot imitate the predictions of much larger teachers. Hence, the KD loss only adds noise to the training signal. At most, on UC2, KD improves the MAE by 1.6$\times$ on light vehicles, at $(e_{dim},d_{dim})=$(96, 64), and by 1.15$\times$ on heavy vehicles, at $(e_{dim},d_{dim})=$(384, 256), compared to regular fine-tuning. In the latter case, the KD-enhanced model reaches a MAEthat is just 1.2$\times$ higher than the teacher while being 3.9$\times$ smaller.
The most interesting results are obtained on UC3. Here, we first note that reducing the model size to $(e_{dim},d_{dim})=$(384-256) is beneficial for performance (5% reduction for a 3.9$\times$ smaller size) even without KD. Moreover, we can achieve very competitive results even with smaller models thanks to our KD-enhanced fine-tuning setup, which is very effective. The most interesting result is achieved at $(e_{dim},d_{dim})=(48,32)$, which further reduces the MAEby 8.6% compared to the baseline, with a model whose parameters only require $0.5$MB, i.e., 195×195$\times$195 × smaller than the original one.
These results show the effectiveness of KD in reducing model size while preserving as much as possible the performance of foundation models from the perspective of real-world deployment resource-constrained devices. The two smallest student configurations require less than 0.7 MBs and could be deployed on the data collection nodes mentioned in Section 4, which are equipped with 1 MB of Flash memory. However, it must be underlined that, on UC2, these downsized models remain hindered by performance gaps that might be too large dependingwon application-specific requirements. Further research is necessary to bridge these gaps, but we believe that our results show promise in the direction of achieving tiny embodied foundation models for SHM.
@@ 7 Conclusions
We proposed a task-independent foundation model for SHM applications, which can achieve state-of-the-art performance on two tasks, Anomaly Detection and Traffic Load Estimation, considering three different datasets. Our transformer-based masked autoencoder learns powerful representations of vibration data gathered by accelerometers during the pre-training step and can be then effectively fine-tuned on the specific downstream task. We believe that our work points to a new and exciting direction in ML for SHM research. Future works may include using a larger and more diverse pre-training dataset, as our results already suggested this could lead to possible benefits, as well as experimenting more thoroughly with model hyper-parameters.

### v2

@@ 6.5 Impact of pre-training
[…]
@@ 6.6 Model size optimization and Knowledge Distillation
Figure 9: MAE% comparison between regular and KD-enhanced fine-tuning. All graphs share the same x-axis.
[…]
These results show the effectiveness of KD in reducing model size while preserving as much as possible the performance of foundation models from the perspective of real-world deployment resource-constrained devices. The two smallest student configurations require less than 0.7 MBs and could be even deployed on the data collection nodes mentioned in Section 4, which are equipped with 1 MB of Flash memory. However, it must be underlined that, on UC2, these downsized models remain hindered by performance gaps that might be too large depending on application-specific requirements.
6.7 Deployment results
Table 6: Deployment results of for a single inference step on NVIDIA Jetson Nano. UC2/UC3 results consider a 60 s input window.
Task UC1 UC2/UC3
Model PCA Ours LR RF k-NN MLP SVR TCN LSTM Ours
Ours (distilled)
Params << 100k 36 M << 100k << 100k << 100k << 100k << 100k 11 M 42 M 25 M 130k
Latency [ms] 14.81 169.36 0.26 0.72 4.63 0.66 0.52 127.91 1533.26 164.93 40.12
While further research is necessary to port these models on SHM sensor nodes, in this section we show that running them in real-time on an edge gateway (present in virtually every sensor node installation, including the ones we considered) is feasible without any loss in accuracy. Table 6 summarizes the deployment results of our architectures on an NVIDIA Jetson Nano, as a proxy for a typical AI-oriented SHM gateway. The module features an NVIDIA Maxwell architecture with 128 CUDA cores and a Quad-core ARM Cortex-A57 MPCore processor, equipped with 4 GB of 64-bit LPDDR4 memory and 16 GB of storage. Inference times were measured across all architectures to assess their real-time feasibility in deployment scenarios, averaging results across 1000 test samples using PyTorch 1.1.0, with model and data in full precision (fp32); because the deployed models are identical to those evaluated offline, their accuracy remains unchanged.
Among all the use cases, the biggest transformer model is the one for UC1 (which includes the Decoder), with 36M parameters: in fp32, this model occupies a total of 144 MB, easily fitting the target gateway’s memory. This model achieves an inference latency of 169.36 ms. While the PCA-based anomaly detector is faster, our transformer is still well below the 2-second constraint imposed by the stride between successive input windows.
For UC2 and UC3, both using the same model where the Decoder has been replaced by a fully connected layer, we achieve 164.93 ms latency, well below the real time constraints of 2 / 15 seconds of UC2 / UC3. As expected, the distilled model is significantly faster, requiring only 40 ms. Classical ML baselines exhibit latencies below 1 ms, which are however unnecessary in this context.
@@ 7 Conclusions
We proposed a task-independent foundation model for SHM applications, which can achieve state-of-the-art performance on two tasks, Anomaly Detection and Traffic Load Estimation, considering three different datasets. Our transformer-based masked autoencoder learns powerful representations of vibration data gathered by accelerometers during the pre-training step and can be then effectively fine-tuned on the specific downstream task. Even if limited to a few SHM tasks, we believe that our work points to a new and exciting direction in ML for SHM research. Future works will include using larger and more diverse pre-training datasets, as our results already suggested this could lead to possible benefits, investigating more SHM tasks involving different civil infrastructures, as well as experimenting more thoroughly with model hyper-parameters.